ben said to his father "in 5 years time you will be twice my age" his dad said "yeah and in 13 yrs time the sum of our ages will be 100" how old is ben now?
needs to be answered using simultaneous equations

Answer by flame8855(424)

y + 5 = 2(x+5)
x+13 +y +13 = 100
x+y = 100- 26
x+y= 74
y = 2x+5
x + 2x +5 =74
3x= 69
x = 23
y = 51